In the majority of night blind parents, the cause is a nutritional deficiency. However, with 8% of cases, the cause is due to a lack of a necessary enzyme to make rhodopsin - a chemical necessary for the rods of the eyes to function. if two parents-to-be have normal vision but are carriers of the gene that can cause the enzyme deficiency how likely is it that their first child will be night blind? If the couple decides to have two children how likely is it that both children will be night-blind?
Answer #1

Let the normal enzyme be A, and the absence of the necessary enzyme be a.

Parents are carriers for this deficiency, hence their genotype is Aa.

Gametes produced by the parents are: A and a.

Parents A a
A AA (normal) Aa (normal, carrier)
a Aa (normal, carrier) aa (night blind)

Thus 1 out of 4 child can be night blind. Probability = 1/4 = 1/4 x 100% = 25%

Probability of the first child being night blind is 1/4th or 25%.

Notwithstanding the phenotype of the first child, the second child also has a 1/4th probability to be night blind.

Probability of both the children being night blind is = 1/4 x 1/4 = 1/16 = 1/16 x 100% = 6.25%.